### v2

# A Sizable Discrepancy in Ground-Based JAGB Distances to Nearby Galaxies

Abstract
Recently, Freedman et al. (2024) report agreement of distances derived from the Tip of the Red Giant Branch (TRGB) and the J-Region Asymptotic Giant Branch (JAGB) at the 1 level for both nearby galaxies with ground-based imaging (0.54 Mpc) as well distant galaxies with JWST imaging (723 Mpc). Here we compare the same ground-based JAGB distances to uniformly reduced space-based optical TRGB distances from the Hubble Space Telescope (HST). We uncover a significant offset between these two distance scales of = 0.17 0.04 (stat) 0.06 (sys) mag (9 in distance), with the HST TRGB distances being further. Inspections of the HST color-magnitude diagrams make a compelling case that the issue lies in the underlying JAGB distances. The source of the disagreement may lie with the lower resolution or photometric calibration of the ground-based near-infrared data, a contrast to the general agreement found between JWST JAGB and other space-based, second-rung distance indicators (Cepheids, Miras, TRGB) presented within Riess et al. (2024). High-resolution, near-infrared observations from an ongoing HST program will enable the simultaneous measurement of Cepheid, JAGB, and TRGB distances in four of these nearby galaxies and allow us to investigate whether the discrepancy noted here is due to ground-based observational systematics, or something intrinsic to the JAGB method relevant for this particular sample. A resolution of this discrepancy is required if the JAGB is to be used to determine a highly precise local value of the Hubble constant.
1 Introduction
Freedman et al. (2024) have recently claimed that the tip of the red giant branch (TRGB; Lee et al. 1993; Madore & Freedman 1995; Beaton et al. 2018; Anand et al. 2021) and J-region asymptotic giant branch (JAGB; Weinberg & Nikolaev 2001; Madore & Freedman 2020; Li et al. 2024) distance scales are in excellent agreement (1), both locally (0.54 Mpc; Freedman & Madore 2020; Lee et al. 2024c), and in more distant galaxies observed with JWST (723 Mpc; Lee et al. 2024a, b). Here we compare ground-based JAGB distances from Lee et al. (2024c) to space-based optical TRGB measurements from the Extragalactic Distance Database’s (EDD) CMDs/TRGB catalog111Available at edd.ifa.hawaii.edu. (Tully et al., 2009; Jacobs et al., 2009; Anand et al., 2021). A key benefit of this limited comparison is that both sets of photometry and distances are publicly available, and measured with the same observatory within each dataset (the JAGB imaging was taken with Magellan-FourStar, whereas the TRGB imaging was taken with the Hubble Space Telescope).
2 Comparison of Distance Scales
The comparison between the two samples is shown in the top-half of Figure 1. The plotted uncertainties in the JAGB distances are re-tabulated from Table 3 in Lee et al. (2024c) to include contributions from the 2MASS photometric zeropoint and 2MASS-to-FourStar correction, error on the JAGB mode, choice of smoothing scale, and a 15 uncertainty in the foreground reddening (). We also include JAGB results for WLM (Lee et al., 2021) and M33 (Lee et al., 2022), which are part of the same ground-based observing program, for a total of thirteen galaxies for comparison. The uncertainties in the TRGB distances are re-tabulated from the CMDs/TRGB catalog to include the statistical uncertainty in the tip measurement and the same 15 uncertainty in foreground reddening (). The mean difference between the two sets of distances is = 0.17 0.04 (stat) 0.06 (sys) mag. The statistical uncertainty is given by , and the systematic uncertainty is given by the quadrature sum of the underlying calibration uncertainty for each method (0.04 mag for each; Rizzi et al. 2007a; Freedman 2021; Lee et al. 2024c). Separately, the blue points show a comparison between the EDD TRGB distances and the ground-based JAGB distances from Zgirski et al. (2021) using data from the InfraRed Survey Facility (IRSF), which show a somewhat increased discrepancy of = 0.23 0.05 (stat) 0.06 (sys) mag, but with only six overlapping galaxies.
The bottom-half of Figure 1 shows four example HST CMDs used for the TRGB measurements. On each of these CMDs, we plot the measured magnitude of the TRGB (in black), and the implied magnitude of the TRGB given by the JAGB distances (in red and blue). For NGC 6822, we see excellent agreement between the TRGB measurement from Anand et al. (2021) and the JAGB value from Lee et al. (2024c). However, for the majority (3/4 cases shown, 9/13 cases total), we see that the JAGB-implied values are inconsistent with the visible onset of the TRGB in these nearby galaxies. While only four examples are shown here, the remainder of the HST CMDs are available via EDD.
3 Potential Solutions
The underlying zeropoint and color-calibration used by the EDD CMDs/TRGB catalog is provided by Rizzi et al. (2007b), with a baseline value of = mag. The average zeropoint for the galaxies in question after including the color calibration is = mag, in good agreement with prior results from the Carnegie-Chicago Hubble Program who find = mag (Freedman, 2021). Given the level of agreement, it seems unlikely that TRGB zeropoint is responsible for the discrepancy.
Could the HST photometry or measurements contained within EDD be subject to large, systematic offsets? Fortunately, subsamples of these 13 galaxies have HST TRGB measurements from other groups, with distinct photometry and TRGB measurement pipelines. The ANGST team (Dalcanton et al., 2009) measured distances to six of the thirteen galaxies compared here, as was also noted by Lee et al. (2024c). Using the ANGST TRGB values collated by Lee et al. (2024c), the average distance between the JAGB and TRGB distance moduli is = 0.21 0.04 (stat) mag, very similar to the offset found in the full comparison using the EDD TRGB sample. Similarly, the GHOSTS team (Radburn-Smith et al., 2011) has measured a TRGB distance to NGC 7793 in three separate fields, and find a mean distance modulus of = 27.87 mag, resulting in an even larger discrepancy between TRGB and JAGB distance scales ( = 0.41 mag) than found with the EDD TRGB measurement ( = 0.34 mag). In short, the HST TRGB distances from the separate groups are all consistently offset in comparison to this sample of JAGB distances to nearby galaxies.
Lee et al. (2024c) mention crowding as a potential cause of concern for NGC 1313, the furthest target in their sample (D = 4.3 Mpc). However, we see large offsets in derived distances beginning at just over 1 Mpc. Fortuitously, an ongoing HST observing program (GO–17520; Breuval et al. 2023) is obtaining high-resolution imaging of four of the galaxies in this sample (NGC 55, NGC 300, NGC 3109, NGC 6822) for the express purpose of performing a 1 cross calibration of the Cepheid, JAGB, and TRGB distance indicators. A WFC3/IR postage stamp of a JAGB star in the outer disk of NGC 55 from this program is shown in the lower-right hand corner of the top panel of Figure 1. A green circle with 1′′ radius is overplotted, and suggests that the ground-based photometry may suffer from the effects of stellar crowding, even in the outer disk regions utilized by Lee et al. (2024c).
In addition to potential crowding, we find that there may be an issue with the photometric calibration of the ground-based data. Lee et al. (2024c) also present near-infrared TRGB distances to many of the same galaxies with the same set of photometry. For eleven galaxies in common, we find a difference of = 0.12 0.04 (stat) 0.07 (sys) mag between our optical HST TRGB and their ground-based NIR-TRGB distances. This disagreement is in the same direction as the disagreement between the HST TRGB and the ground-based JAGB measurements, albeit at a somewhat reduced level (and with less significance). Given that the TRGB stars used by Lee et al. (2024c) lie in significantly less dense regions than their JAGB stars and are thus less susceptible to crowding, the underlying photometric calibration may also be a culprit.
It seems likely that crowding and/or photometric calibration are a significant part of the solution, but it is worth pointing out that the nearby galaxies in the comparison presented here are somewhat lower mass than those used in the Cepheid+Type Ia distance ladder (Riess et al., 2022). For those galaxies, there is good general agreement between JWST JAGB measurements and those from other second-rung distance indicators (Li et al., 2024; Lee et al., 2024b; Riess et al., 2024). Notably, JWST imaging in the 7-23 Mpc range is of higher resolution than ground-based imaging in the 2-4 Mpc range. Future investigations into the dependence of the JAGB on metallicities, ages, and full star-formation histories are duly warranted (Lee, 2023; Li et al., 2024).